# Indirect forces in disc-planet interaction

### v1

@@ 5.2 High mass planets
The situation is different for more massive planets that can carve deep gaps around their orbits. Gap opening strongly suppresses the direct contribution of $m\sim h_{\mathrm{p}}^{-1}$ Lindblad resonances to $T_{\mathrm{p}}$ as $\Sigma$ is dramatically lowered at their locations. In this case the indirect Lindblad torque contribution $T_{\mathrm{L,id}}$ at the $m=1$ OLR would grow in relative significance (compared to $T_{\mathrm{p,d}}$) as its location $R_{1}=2^{2/3}R_{\mathrm{p}}$ is far from $R_{\mathrm{p}}$ and $\Sigma$ there will not be suppressed nearly as much as around $R_{\mathrm{p}}$. In the most extreme case, when $\Sigma$ is strongly reduced everywhere except the $m=1$ OLR, only $T_{\mathrm{L,id}}$ and the $m=1$ direct Lindblad contribution $T_{\mathrm{L,d}}$ would contribute to the torque on the planet $T_{\mathrm{p}}$.
As we show in Appendices C.2 and C.3, these two torque contributions are comparable in magnitude but different in sign, with $T_{\mathrm{L,id}}\approx-0.74\,T_{\mathrm{L,d}}$ (see equations (50), (51), Appendices C.2-C.4 and Figure 13). As a result, the full torque on the planet coming from the $m=1$ OLR $T_{\mathrm{p}}=T_{\mathrm{L,d}}+T_{\mathrm{L,id}}$ is roughly 4 times smaller in magnitude than just the direct component $T_{\mathrm{L,d}}$ This implies that in studies of disc-planet coupling in high-$M_{\mathrm{p}}$ regime, the indirect torque of the disc on the planet can be very significant and should be explicitly taken into account in the calculation of $T_{\mathrm{p}}$. This is definitely necessary when the planet is capable of opening a gap or cavity around its orbit and its migration switches to the Type II regime (Ward1997). An analogous conclusion was reached by Crida2025 through direct numerical simulations.
@@ 6 Discussion
The main goal of our study was to understand the quantitative impact of excluding the IT from the calculation of disc-planet interaction in a non-inertial, astrocentric frame. Note that we are not questioning the conceptual necessity of including the IT in this frame — its presence unambiguously follows from the equations of motion in the non-inertial frame (see Section 2.1). We are in agreement with Crida2025 that the IT associated with a particular component of the system (e.g. a planet or a disc) should be included/excluded when the direct gravitational effect of that component is included/excluded in the calculation. But here we are interested in the actual quantitative differences arising when the IT is dropped from the calculation.
@@
Somewhat ironically, the same indirect torque oscillations also reflect the non-conservation of the angular momentum of the full star-planet-disc system in the non-inertial astrocentric frame, which is evident from equation (33), keeping in mind that the stellar angular momentum is zero in this frame. This non-conservation arises as a result of the torque exerted by the indirect force $M\textbf{{a}}_{\mathrm{id}}$ ($M$ is the total mass of the system including $M_{\star}$) applied to the barycenter of the system, as shown in Rafikov2025.
@@
We would like to re-iterate that the indirect torque oscillations should not be confused with the torque wiggles studied in Cimerman2024b or the torque oscillations in circumbinary discs explored in Cimerman2024a. These phenomena arise because of the direct gravitational force of the planet coupling with the global planet-driven spiral density wave on scales $\gtrsim R_{\mathrm{p}}$. On the other hand, the indirect torque oscillations are due to the indirect force coupling to the same density wave on global scales. The difference is easily seen in the behavior of the oscillation amplitudes — the indirect torque oscillations grow with $R$, whereas torque wiggles decay with $R$ (Cimerman2024b). The two types of torque oscillations actually co-exist, as can be noticed in Figures 4-7 (the first torque wiggle around $2.2R_{\mathrm{p}}$ can be seen in both $\mathrm{d}T/\mathrm{d}R$ and $F_{J}$ curves).
@@
Despite the divergent oscillations at large $R$ exhibited by both the wave AMF and the full integrated torque $T=T_{\mathrm{d}}+T_{\mathrm{id}}$, the deposition torque density $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ — the difference between the two — is well behaved and decays to zero as $R\to\infty$ (inasmuch as numerical artefacts allow one to determine it), see Figure 8. Moreover, the profile of $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ is largely independent of whether it is computed in the non-inertial frame, barycentric frame, or not accounting for the IT at all, at least for low (sub-$M_{\mathrm{th}}$) mass planets. This is good news for studies of gap opening by planets, since the behavior of $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ — the amount of angular momentum absorbed by the disc material per unit radius — is the key driver of disc evolution resulting from planet-disc coupling. Thus, even studies that disregard the IT altogether should still provide a reliable description of gap opening by the low mass planets. However, the situation may change in the high-$M_{\mathrm{p}}$ regime, when a planet opens a deep gap, if the direct excitation torque density $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$ ends up being very different between the calculations including and disregarding the IT due to the planet.
@@
Similarly, the amount of angular momentum absorbed (or lost) by a low mass (sub-$M_{\mathrm{th}}$) planet changes weakly, at the level of several per cent, if the IT due to the disc is turned off, as we demonstrated in Section 5.1. Such small changes to the torque on the planet $T_{\mathrm{p}}$ are expected to only slightly alter the planetary migration rate in the Type I regime. In this study we considered only circular planets, but previously Fairbairn2025a showed that even for a non-circular planet, its eccentricity evolution does not change dramatically when the IT is switched off since $T_{\mathrm{p}}$ changes very little. However, for high mass planets, capable of opening deep gaps around their orbit, the situation may be very different, as the contribution to the torque coming from the normally dominant high $m\sim h_{\mathrm{p}}^{-1}$ Lindblad resonances close to the planet (GT80) gets suppressed. As a result, the indirect component of the disc torque on the planet excited at $m=1$ OLR starts playing a very important role (see Section 5.2). Thus, the indirect component of this torque may significantly affect the rate of Type II migration for massive planets.
@@
Summarizing this discussion, we find that neglecting the IT in disc-planet interaction calculations often results in only small differences that can be tolerated. In particular, for low mass planets, in the linear regime, neither the disc evolution (i.e. gap opening) nor the orbital evolution of the planet (its radial migration) should change by much. Nevertheless, fully accounting for the IT due to some massive component of the system is still highly recommended whenever the direct gravitational effect of that component is included, even if just for self-consistency (Crida2025). And in some cases, e.g. for higher mass planets, the IT should play an important quantitative role in evolution of the system (e.g. see Section 5.2).
@@ 6.1 Boundary conditions for simulations
In reality, far from the planet ($R\gg R_{\mathrm{p}}$) the disc tends to orbit around the system barycenter, so the natural axisymmetric BCs should be applied at fixed $r$. Applying them at fixed $R$, as routinely done in astrocentric simulations, leads to azimuthally-varying deviations of the fluid velocities at the outer boundary relative to the true BCs, the amplitude of which scales $\propto M_{\mathrm{p}}/\mathcal{M}_{\star}$. As a result, a parasitic perturbation is forced in the flow at the outer boundary which may propagate through the whole domain and contaminate the results of a simulation. This problem has been previously noted [^9] Earlier, KP93 noted that ‘The indirect term contributes not only in the body of the disk, but also through a term that is strongly dependent in a complicated way on the exact treatment of the disk edge.’ in Teys2017 and a recipe for dealing with it was outlined in Dempsey2021. We independently rediscovered this issue and recommend mitigating it in simulations by applying the appropriate azimuthally-varying outer BCs at fixed $R$.
@@ 6.2 Comparison with existing literature
There is an extensive body of literature on disc-planet coupling, both numerical and analytical. Some of these studies include the IT, while others do not; we do not name specific examples as the lists would be too long both ways. In some cases the inclusion or neglect of the IT is clearly specified, but often it is not mentioned explicitly. We support Crida2025 in urging everyone to explicitly state whether the IT due to each component of the system under study is included (or not) in the calculations performed in the astrocentric frame.
@@
When the IT is explicitly neglected, this approach is usually justified on the grounds of simplicity, or because it does not make large differences in the outcomes, although typically no quantitative assessment of these differences is provided. KP93 and Miranda2019I also motivated the neglect of the IT for practical purposes, since its inclusion raises compatibility issues with the radiative boundary conditions adopted in their linear calculations. As our results demonstrate, in many cases (predominantly for low-mass planets) the neglect of the IT leads to only minor changes in the evolution of the disc and planetary orbit, corroborating the results of earlier studies not accounting for the IT.
@@
We are not aware of any studies that tried to explicitly compare the perturbation patterns induced by a planet with or without the planetary IT, as we do in Section 3. The planetary torque on the disc has been previously computed by a number of authors in different ways. For example, Bate2003, Duffell2012, Cimerman2024b do not include the planetary IT in their calculations of the $\Sigma$ perturbation, while Kley2012, Arza18 do. However, in all of these cases, only the direct torque component was computed, so that $T=T_{\mathrm{d}}$ was assumed. To the best of our knowledge, prior to our work no other study of planet-disc coupling accounted for the indirect torque. It was considered only in the context of modeling discs in stellar binaries (cataclysmic variables) by Ju2016; Ju2017. However, since such discs lie interior to the orbit of the perturber, the indirect torque plays a negligible role (in the absence of permanent eccentricity) and exhibits no prominent oscillations (see Section 4.4).
@@
Similarly, not many planet-disc studies computed the AMF $F_{J}$ in a global setting, and when this was done, the IT was neglected altogether (e.g. Cimerman2021, Cimerman2024b). The work of Zia2023 represents an interesting exception: they include the planetary IT in the calculation of the disc perturbation in the astrocentric frame and also compute $F_{J}$. They comment on ‘some oscillations in the radial profiles of the AMF’ in the outer disc (see their Fig. 5), which, in light of our results, should be interpreted as the AMF oscillations reflecting the effect of the indirect torque $T_{\mathrm{id}}$. Brown2025 also accounted for the planetary IT in the calculation of the $\Sigma$ perturbation and computed $F_{J}$ but only in the disc interior to the planetary orbit, where no interesting effects due to the indirect forces are expected.
@@
Studies of planetary orbital evolution typically neglected indirect forces (KP93; Pap2000). The situation changed recently when Fairbairn2025a computed planetary migration and eccentricity damping timescales, both with and without the disc IT. Working in the linear regime ($M_{\mathrm{p}}\ll M_{\mathrm{th}}$), they found these timescales to change at most by $\sim 10\%$ when the disc IT was not included. Later, Crida2025 supported this conclusion with direct numerical simulations. They also found that the disc IT plays a more substantial role in planet migration for high $M_{\mathrm{p}}$, in the Type II migration regime. All of these findings are in full agreement with our (semi-)analytical results in Section 5 and Appendix C.
@@ Acknowledgements
We are grateful to Lev Arzamasskiy, Alexandros Ziampras, Scott Tremaine and Jim Stone for useful discussions. R.R.R. acknowledges financial support from the STFC grant ST/T00049X/1 and the IBM Einstein Fellowship at the IAS. A.J.D. was supported by NASA through the NASA Hubble Fellowship grant #HST-HF2-51553.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. C.W.F acknowledges support provided by the Friends of the IAS Members fund.
@@
Here we explain the non-monotonic behavior of $\Sigma_{m}$ featuring multiple nulls in Figure 3, illustrating a mildly non-linear case with $M_{\mathrm{p}}=0.3M_{\mathrm{th}}$. We show that this behavior has its origin in the azimuthal broadening of the planet-driven density wave caused by nonlinear effects (GR01), which is evident in Figure 3a.
@@ Appendix B Torque density expressions
(see equation (A4) of Cimerman2024a) with
@@
being the Laplace coefficients (e.g. MurrayDermott), equation (44) becomes
@@
Here we introduced an indicator exponent $\zeta=\Theta(R_{\mathrm{p}}-R)$ with $\Theta(z)$ being the Heavyside step-function, making the expression (47) applicable in both the inner ($R<R_{\mathrm{p}}$, $\zeta=1$) and the outer ($R>R_{\mathrm{p}}$, $\zeta=0$) parts of the disc. In a somewhat different form this expression has been previously derived in Cimerman2024b; its generalization in an inertial, barycentric frame has been obtained in Cimerman2024a in the context of circumbinary discs.

### v3

@@ 5.2 High mass planets
The situation is different for more massive planets that can carve deep gaps around their orbits. Gap opening strongly suppresses the direct contribution of $m\gtrsim h_{\mathrm{p}}^{-1}$ Lindblad resonances to $T_{\mathrm{p}}$ as $\Sigma$ is dramatically lowered at their locations. In this case the indirect Lindblad torque contribution $T_{\mathrm{L,id}}$ at the $m=1$ OLR would grow in relative significance (compared to $T_{\mathrm{p,d}}$) as its location $R_{1}=2^{2/3}R_{\mathrm{p}}$ is far from $R_{\mathrm{p}}$ and $\Sigma$ there will not be suppressed nearly as much as around $R_{\mathrm{p}}$. In the most extreme case, when $\Sigma$ is strongly reduced everywhere except the $m=1$ OLR, only $T_{\mathrm{L,id}}$ and the $m=1$ direct Lindblad contribution $T_{\mathrm{L,d}}$ would contribute to the torque on the planet $T_{\mathrm{p}}$.
As we show in Appendices C.2 and C.3, these two torque contributions are comparable in magnitude but different in sign, with $T_{\mathrm{L,id}}\approx-0.74\,T_{\mathrm{L,d}}$ (see equations (50), (51), Appendices C.2-C.4 and Figure 14). As a result, the full torque on the planet coming from the $m=1$ OLR $T_{\mathrm{p}}=T_{\mathrm{L,d}}+T_{\mathrm{L,id}}$ is roughly 4 times smaller in magnitude than just the direct component $T_{\mathrm{L,d}}$ This implies that in studies of disc-planet coupling in high-$M_{\mathrm{p}}$ regime, the indirect torque of the disc on the planet can be very significant and should be explicitly taken into account in the calculation of $T_{\mathrm{p}}$. This is definitely necessary when the planet is capable of opening a gap or cavity around its orbit and its migration switches to the Type II regime (Ward, 1997). An analogous conclusion was reached by Regály and Vorobyov (2017) and Crida et al. (2025) through direct numerical simulations.
@@ 6 Discussion
The main goal of our study was to understand the quantitative impact of excluding the IT from the calculation of disc-planet interaction in a non-inertial, astrocentric frame. Note that we are not questioning the conceptual necessity of including the IT in this frame — its presence unambiguously follows from the equations of motion in the non-inertial frame (see Section 2.1). We are in agreement with Crida et al. (2025) that the IT associated with a particular component of the system (e.g. a planet or a disc) should be included/excluded when the direct gravitational effect of that component is included/excluded in the calculation. But here we are interested in the actual quantitative differences arising when the IT is dropped from the calculation.
@@
Somewhat ironically, the same indirect torque oscillations also reflect the non-conservation of the angular momentum of the full star-planet-disc system in the non-inertial astrocentric frame, which is evident from equation (33), keeping in mind that the stellar angular momentum is zero in this frame. This non-conservation arises as a result of the torque exerted by the indirect force $M\textbf{{a}}_{\mathrm{id}}$ ($M$ is the total mass of the system including $M_{\star}$) applied to the barycenter of the system, as shown in Rafikov (2025).
@@
We would like to re-iterate that the indirect torque oscillations should not be confused with the torque wiggles studied in Cimerman et al. (2024) or the torque oscillations in circumbinary discs explored in Cimerman and Rafikov (2024). These phenomena arise because of the direct gravitational force of the planet coupling with the global planet-driven spiral density wave on scales $\gtrsim R_{\mathrm{p}}$. On the other hand, the indirect torque oscillations are due to the indirect force coupling to the same density wave on global scales. The difference is easily seen in the behavior of the oscillation amplitudes — the indirect torque oscillations grow with $R$, whereas torque wiggles decay with $R$ (Cimerman et al., 2024). The two types of torque oscillations actually co-exist, as can be noticed in Figures 4-7 (the first torque wiggle around $2.2R_{\mathrm{p}}$ can be seen in both $\mathrm{d}T/\mathrm{d}R$ and $F_{J}$ curves).
@@
Despite the divergent oscillations at large $R$ exhibited by both the wave AMF and the full integrated torque $T=T_{\mathrm{d}}+T_{\mathrm{id}}$, the deposition torque density $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ — the difference between the two — is well behaved and decays to zero as $R\to\infty$ (inasmuch as numerical artefacts allow one to determine it), see Figure 9. Moreover, the profile of $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ is largely independent of whether it is computed in the non-inertial frame, barycentric frame, or not accounting for the IT at all, at least for low (sub-$M_{\mathrm{th}}$) mass planets. This is good news for studies of gap opening by planets, since the behavior of $\mathrm{d}T_{\mathrm{dep}}/\mathrm{d}R$ — the amount of angular momentum absorbed by the disc material per unit radius — is the key driver of disc evolution resulting from planet-disc coupling. Thus, even studies that disregard the IT altogether should still provide a reliable description of gap opening by the low mass planets.
However, the situation should change in the high-$M_{\mathrm{p}}$ regime, when a planet opens a deep gap, if the direct excitation torque density $\mathrm{d}T_{\mathrm{d}}/\mathrm{d}R$ ends up being very different between the calculations including and disregarding the IT due to the planet. In particular, Regály and Vorobyov (2017) found that inclusion of the IT in their simulations decreases the width of the gap carved by a planet. This may be a consequence of the lower amplitude of $\delta\Sigma_{1}$ (see Figures 2c & 3c) at $m=1$ OLR when the IT is included, weakening disc-planet coupling and reducing both the total direct planetary torque exerted on the outer disc and the deposited torque.
@@
Similarly, the amount of angular momentum absorbed (or lost) by a low-mass (sub-$M_{\mathrm{th}}$) planet changes weakly, at the level of several per cent, if the IT due to the disc is turned off, as we demonstrated in Section 5.1. Such small changes to the torque on the planet $T_{\mathrm{p}}$ are expected to only slightly alter the planetary migration rate in the Type I regime. In this study we considered only circular planets, but previously Fairbairn and Rafikov (2025) showed that even for a non-circular planet, its eccentricity evolution does not change dramatically when the IT is switched off since $T_{\mathrm{p}}$ changes very little. However, for high mass planets, capable of opening deep gaps around their orbit, the situation may be very different, as the contribution to the torque coming from the normally dominant high $m\sim h_{\mathrm{p}}^{-1}$ Lindblad resonances close to the planet (Goldreich and Tremaine, 1980) gets suppressed. As a result, the indirect component of the disc torque on the planet excited at $m=1$ OLR starts playing a very important role (see Section 5.2). Thus, the indirect component of this torque may significantly affect the rate of Type II migration for massive planets.
@@
Summarizing this discussion, we find that neglecting the IT in disc-planet interaction calculations often results in only small differences that can be tolerated. In particular, for low mass planets, in the linear regime, neither the disc evolution (i.e. gap opening) nor the orbital evolution of the planet (its radial migration) should change by much. Nevertheless, fully accounting for the IT due to some massive component of the system is still highly recommended whenever the direct gravitational effect of that component is included, even if just for self-consistency (Crida et al., 2025). And in some cases, e.g. for higher mass planets, the IT should play an important quantitative role in evolution of the system (e.g. see Section 5.2).
@@ 6.1 Boundary conditions for simulations
In reality, far from the planet ($R\gg R_{\mathrm{p}}$) the disc tends to orbit around the system barycenter, so the natural axisymmetric BCs should be applied at fixed $r$. Applying them at fixed $R$, as routinely done in astrocentric simulations, leads to azimuthally-varying deviations of the fluid velocities at the outer boundary relative to the true BCs, the amplitude of which scales $\propto M_{\mathrm{p}}/\mathcal{M}_{\star}$. As a result, a parasitic perturbation is forced in the flow at the outer boundary which may propagate through the whole domain and contaminate the results of a simulation. This problem has been previously noted [^9] Earlier, Korycansky and Pollack (1993) noted that ‘The indirect term contributes not only in the body of the disk, but also through a term that is strongly dependent in a complicated way on the exact treatment of the disk edge.’ in Teyssandier and Ogilvie (2017) and a recipe for dealing with it was outlined in Dempsey et al. (2021, see their Appendix B). We independently rediscovered this issue and recommend mitigating it in simulations by applying the appropriate azimuthally-varying outer BCs at fixed $R$.
@@ 6.2 Comparison with existing literature
There is an extensive body of literature on disc-planet coupling, both numerical and analytical. Some of these studies include the IT, while others do not; we do not name specific examples as the lists would be too long both ways. In some cases the inclusion or neglect of the IT is clearly specified, but often it is not mentioned explicitly. We support Crida et al. (2025) in urging everyone to explicitly state whether the IT due to each component of the system under study is included (or not) in the calculations performed in the astrocentric frame.
@@
When the IT is explicitly neglected, this approach is usually justified on the grounds of simplicity, or because it does not make large differences in the outcomes, although typically no quantitative assessment of these differences is provided. Korycansky and Pollack (1993) and Miranda and Rafikov (2019a) also motivated the neglect of the IT for practical purposes, since its inclusion raises compatibility issues with the radiative boundary conditions adopted in their linear calculations. As our results demonstrate, in many cases (predominantly for low-mass planets) the neglect of the IT leads to only minor changes in the evolution of the disc and planetary orbit, corroborating the results of earlier studies not accounting for the IT.
@@
We are not aware of any studies that tried to explicitly compare the perturbation patterns induced by a planet with or without the planetary IT, as we do in Section 3. The planetary torque on the disc has been previously computed by a number of authors in different ways. For example, Bate et al. (2003), Duffell and MacFadyen (2012), Cimerman et al. (2024) do not include the planetary IT in their calculations of the $\Sigma$ perturbation, while Kley et al. (2012), Arzamasskiy et al. (2018) do. However, in all of these cases, only the direct torque component was computed, so that $T=T_{\mathrm{d}}$ was assumed. To the best of our knowledge, prior to our work no other study of planet-disc coupling accounted for the indirect torque. It was considered only in the context of modeling discs in stellar binaries (cataclysmic variables) by Ju et al. (2016, 2017). However, since such discs lie interior to the orbit of the perturber, the indirect torque plays a negligible role (in the absence of permanent eccentricity) and exhibits no prominent oscillations (see Section 4.4).
@@
Similarly, not many planet-disc studies computed the AMF $F_{J}$ in a global setting, and when this was done, the IT was neglected altogether (e.g. Cimerman and Rafikov 2021, Cimerman et al. 2024). The work of Ziampras et al. (2023) represents an interesting exception: they include the planetary IT in the calculation of the disc perturbation in the astrocentric frame and also compute $F_{J}$. They comment on ‘some oscillations in the radial profiles of the AMF’ in the outer disc (see their Fig. 5), which, in light of our results, should be interpreted as the AMF oscillations reflecting the effect of the indirect torque $T_{\mathrm{id}}$. Brown and Ogilvie (2025) also accounted for the planetary IT in the calculation of the $\Sigma$ perturbation and computed $F_{J}$ but only in the disc interior to the planetary orbit, where no interesting effects due to the indirect forces are expected.
@@
Studies of planetary orbital evolution typically neglected indirect forces (Korycansky and Pollack, 1993; Papaloizou and Larwood, 2000). The situation changed recently when Fairbairn and Rafikov (2025) computed planetary migration and eccentricity damping timescales, both with and without the disc IT. Working in the linear regime ($M_{\mathrm{p}}\ll M_{\mathrm{th}}$), they found these timescales to change at most by $\sim 10\%$ when the disc IT was not included. Later, Crida et al. (2025) supported this conclusion with direct numerical simulations. They also found that the disc IT plays a more substantial role in planet migration for high $M_{\mathrm{p}}$, in the Type II migration [^10] Regály and Vorobyov (2017) also found different Type II migration rates when allowing or prohibiting the star to move in their simulations. However, they ascribe this difference to a different gap structure between the two cases, which is a consequence of the planetary IT rather than the disc IT. regime. All of these findings are in full agreement with our (semi-)analytical results in Section 5 and Appendix C.
@@ Acknowledgements
We are grateful to Lev Arzamasskiy, Alexandros Ziampras, Scott Tremaine and Jim Stone for useful discussions, and to an anonymous referee for helpful suggestions. R.R.R. acknowledges financial support from the STFC grant ST/T00049X/1 and the IBM Einstein Fellowship at the IAS. A.J.D. was supported by NASA through the NASA Hubble Fellowship grant #HST-HF2-51553.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. C.W.F acknowledges support provided by the Friends of the IAS Members fund.
@@
Here we explain the non-monotonic behavior of $\Sigma_{m}$ featuring multiple nulls in Figure 3, illustrating a mildly non-linear case with $M_{\mathrm{p}}=0.3M_{\mathrm{th}}$. We show that this behavior has its origin in the azimuthal broadening of the planet-driven density wave caused by nonlinear effects (Goodman and Rafikov, 2001), which is evident in Figure 3a.
@@ Appendix B Torque density expressions
(see equation (A4) of Cimerman and Rafikov, 2024) with
@@
being the Laplace coefficients (e.g. Murray and Dermott, 1999), equation (44) becomes
@@
Here we introduced an indicator exponent $\zeta=\Theta(R_{\mathrm{p}}-R)$ with $\Theta(z)$ being the Heavyside step-function, making the expression (47) applicable in both the inner ($R<R_{\mathrm{p}}$, $\zeta=1$) and the outer ($R>R_{\mathrm{p}}$, $\zeta=0$) parts of the disc. In a somewhat different form this expression has been previously derived in Cimerman et al. (2024); its generalization in an inertial, barycentric frame has been obtained in Cimerman and Rafikov (2024) in the context of circumbinary discs.